OpenStudy (anonymous):

The radii of Earth and Pluto are 6,371 kilometers and 1,161 kilometers, respectively. Approximately how many spheres the size of Pluto does it take to have the same volume as Earth?

OpenStudy (jdoe0001):

\(\bf \textit{volume of a sphere} = \cfrac{4}{3}\pi r^3\\ \quad \\ Pluto's = \cfrac{4}{3}\cdot\pi\cdot 1161^3\\ \quad \\ Earth's = \cfrac{4}{3}\cdot\pi\cdot 6371^3\)